OpenStudy (anonymous):

parallel to 4x+4y+7=0 and passing through (2,1)

OpenStudy (anonymous):

solve for y it makes it a little easier

OpenStudy (anonymous):

so -4y = 4x+7 y = (4/-4)x +7/4 we can ignore the equations y intecercept because we only care about having teh same slope y = -1x now plug in teh point to find teh new equation y = -1x +b 1 = -1(2) +b 1 = -2 + b 1+2 = b 3 = b y = -1x +3

OpenStudy (ankit042):

find the slope first then use y=mx+c where m is slope

OpenStudy (solomonzelman):

My teacher would give me a zipp if I said `y=-1x+3`

OpenStudy (solomonzelman):

`y=-x+3` ..... my teacher would...

OpenStudy (ankit042):

haha

OpenStudy (anonymous):

@SolomonZelman whats a zipp

OpenStudy (solomonzelman):

I am saying, "zero"

OpenStudy (anonymous):

@SolomonZelman why

OpenStudy (ankit042):

He mean to say that you should not write 1x as it is same as writing x But I feel its fine don't worry